How do you SOlve a linear equation like this one
3(4z-1)-2(2+2)=2(z+1)
And
-9.5x- 0.05= 10.5x + 1.05

Found 2 solutions by queenofit, bacos:
Answer by queenofit(120)

The first thing you must understand is the numbers (or variables)in front of the ( ) are actually being multiplied to the numbers and/or variables inside the ( ). So first lets distribute the numbers that need to be multiplied throught your equation like this. your problem is stated
3(4z-1)-2(2+2)=2(z+1) first I will take the 3 and mult. into the ( ).
12z-3-2(2+2)=2(z+1) I do the rest the same way.
12z-3-4+4=2z+2 do you see how I multiplied the numbers outside () to inside?
then you move your numbers and variables to opposite sides of the = signs, remember to change your +/- signs when you cross the = sign. so you will have
12z-2z=3+4-4+2
10z=5 you still have a 10 in front, so you must divide each side by 10
10z/10=5/10
z=5/10
z=1/2
does this help?

Answer by bacos(9)

Simplifying
3(4z + -1) + -2(2 + 2) = 2(z + 1)
Reorder the terms:
3(-1 + 4z) + -2(2 + 2) = 2(z + 1)
(-1 * 3 + 4z * 3) + -2(2 + 2) = 2(z + 1)
(-3 + 12z) + -2(2 + 2) = 2(z + 1)
Combine like terms: 2 + 2 = 4
-3 + 12z + -2(4) = 2(z + 1)
Multiply -2 * 4
-3 + 12z + -8 = 2(z + 1)
Reorder the terms:
-3 + -8 + 12z = 2(z + 1)
Combine like terms: -3 + -8 = -11
-11 + 12z = 2(z + 1)
Reorder the terms:
-11 + 12z = 2(1 + z)
-11 + 12z = (1 * 2 + z * 2)
-11 + 12z = (2 + 2z)
Solving
-11 + 12z = 2 + 2z
Solving for variable 'z'.
Move all terms containing z to the left, all other terms to the right.
Add '-2z' to each side of the equation.
-11 + 12z + -2z = 2 + 2z + -2z
Combine like terms: 12z + -2z = 10z
-11 + 10z = 2 + 2z + -2z
Combine like terms: 2z + -2z = 0
-11 + 10z = 2 + 0
-11 + 10z = 2
Add '11' to each side of the equation.
-11 + 11 + 10z = 2 + 11
Combine like terms: -11 + 11 = 0
0 + 10z = 2 + 11
10z = 2 + 11
Combine like terms: 2 + 11 = 13
10z = 13
Divide each side by '10'.
z = 1.3
Simplifying
z = 1.3